explain: -14-8+10-(-6)= ??
I get = 10 and 38
help I have trouble only with the substraction. I can do the addition, division and multiplication of positive and negative numbers. I just do not get the substraction of - / + nos.
Thanks

Answer by mystic1306(20)

-14-8+10-(-6)
first -(-6) makes 6 because you are multipling a negitive times a positive which make a positive
rewrite the problem
-14-8+10+6
now take it in steps
-14-8= -22
10+6= 16
-22+16= -6
So..
-14-8+10-(-6)=
-14-8 +10+6=
-22 + 16 = -6